I just need someone to review my work.
Question:
If the cholesterol level of men in the community is normally distributed with a mean of 200 and a standard deviation of 25, what is the probability that a randomly selected sample of of 49 men will have a mean between 190 and 205? (hint: this question pertains to the sampling distribution of the mean).
(190-200)/(25/√49)= -10/3.57= -2.80
(205-200)/(25//√49)= 5/3.57= 1.4
P(X-Bar > 190) = 0.50 –.4974= .0026
0.26% of sample means (from repeated samples of n=49 men) will have cholesterol mean value greater than 190.

P(X-Bar <205) = 0.50 + .4192 = .9192
91.92% of sample means (from repeated samples of n=49) will have a cholesterol level less than 205.


Answer by ewatrrr(24785)

 
Hi,
RE TY, Yes, NORMSDIST(-2.8) is .00256 and NORMSDIST(1.4) is .91924
.91924
-.00256 is .91668 Or = .9167 Or 91.67%
(Ans Correct, as listed, rounded to 4 decimal points, calculation Using 5)
As to: .9192 - .0026 (Using 4 decimal points for calculation...)= 91.66%
t = (190-200)/(25/√49)= -10/3.57= -2.80 Yes
t = (205-200)/(25//√49)= 5/3.57= 1.4 Yes
P(180< x <205) = P( -2.80 ≤ t ≤ 1.4) = .91924 - .00256 = .9167